OpenStudy (anonymous):

2(x + 8) − (x + 1) = 20

OpenStudy (anonymous):

hurry sasoo

OpenStudy (sasogeek):

distribute, put like terms together, solve for x

OpenStudy (sasogeek):

can u do that?

OpenStudy (anonymous):

waitt

OpenStudy (bahrom7893):

2x + 16 - x - 1 = 20, keep all xs on one side and throw everything else on the other.

OpenStudy (anonymous):

i already have ihtt

OpenStudy (sasogeek):

i wonder who the slow one is...

OpenStudy (bahrom7893):

hey.. i never called u slow..

OpenStudy (sasogeek):

her lol :P view our previous thread, u'll know what i mean

OpenStudy (anonymous):

2(x + 8) − (x + 1) = 20 2x + 16 - x -1 = 20 x = 20 - 15 x = 5